OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

Well, what values should we use for h and r?

OpenStudy (anonymous):

Um, im not sure

OpenStudy (anonymous):

h is the height above the ground. we need to find t when the ball hits the ground. What should h be equal to ?

OpenStudy (anonymous):

0 ? i dont know

OpenStudy (anonymous):

Yes, because the ball starts from ground level and ends up there too.

OpenStudy (anonymous):

What about r? That's the speed of the ball.

OpenStudy (anonymous):

96

OpenStudy (anonymous):

Correct. Notice that r > 0 and the coefficient of t^2 is < 0. That's because the ball starts out going up, and gravity accelerates downward. Now all you have to do is solve -16t^2 + 96t = 0

OpenStudy (anonymous):

Why not multiply both sides by 1/16 to start with?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-t^2 + 6t = 0

OpenStudy (anonymous):

Now, one solution to this is 0. That's the solution when the ball is just starting up. If we multiply both sides by 1/t that's OK in this case. What do you get?

OpenStudy (anonymous):

i dont understand that part ?

OpenStudy (anonymous):

(1/t) (-t^2 + 6t) = (1/t) 0 = 0 -t + 6 = 0

OpenStudy (anonymous):

Do you understand that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

t = 6

OpenStudy (anonymous):

Now, plug 6 back into the original equation to make sure it's correct, and you're done.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sorry, posted to the wrong thread, but I hope you are happy with what we did

OpenStudy (anonymous):

you are very welcome